# Large Language Models and Code Security: A Systematic Literature Review

### v2

@@ IX-B Impact of Poisoned Training Data on Ability to Detect and Fix Vulnerabilities
Although research on data poisoning in LLMs exists, there are currently no studies that examine its particular effects on vulnerability detection and fixing tasks. This gap in the literature will be further discussed in the Open challenges, in Section X-B.
X Discussion
@@
In Section III, as a part of our methodology, we have defined three RQs that steered the focus of this study. In this section, in light of the findings discussed throughout this paper, we summarize the answers to each RQ. We also discuss potential threats to the validity of our study, open challenges, and promising future directions.
X-A Addressing Research Questions
@@
The literature revealed that LLM-generated code often contains different security vulnerabilities. These vulnerabilities are largely due to inadequate input validation, improper memory and resource management, weak cryptographic practices, and poor file handling. Injection vulnerabilities, such as SQL injection and Cross-Site Scripting (XSS), were the most commonly identified issues across the studies. Although LLMs can speed up the code generation process, they can produce vulnerable code due to their lack of understanding of secure coding practices.
@@
Studies have demonstrated that LLMs trained on poisoned datasets can suggest insecure code, such as using insecure encryption methods (e.g., ECB). These suggestions can lead programmers to include inadvertently security vulnerabilities in their code. Additionally, studies have proposed different poisoning attack methods that can successfully bypass detection by SATs and LLMs, for example, by hiding malicious snippets in less visible places. Studies have shown that even a small fraction of poisoned training data (i.e., less than 3%) could lead a model to produce vulnerable code. Currently, no research addresses in-depth the effectiveness of poisoned LLM models, when tasked to detecting and fixing vulnerabilities.
X-B Open Challenges
@@
Although LLMs show great potential in code generation and security-related tasks, they still face challenges that need to be addressed. As discussed in Section IV, LLM-generated code often introduces security vulnerabilities. Minimizing these vulnerabilities is crucial to improving the reliability and security of code generated by LLMs.
@@
Besides the aforementioned challenges faced in producing secure code, LLMs encounter challenges also when analyzing code for detecting vulnerabilities in it. As we have uncovered in Subsection V-B, the first and main issue is the one of false positives, which severely limits current LLMs capabilities in vulnerability detection. Reducing false positive rates is necessary for ensuring higher reliability of LLMs and their broader adoption for undertaking these tasks. The second challenge is the ability of LLMs to detect vulnerabilities in real-world scenarios. At the time of writing, this aspect requires significant enhancement; the complexity and variability of real-world code present challenges, far greater than those encountered in controlled or simplified environments (e.g., most vulnerability datasets).
@@
Third, LLMs encounter substantial challenges when it comes to fixing vulnerabilities in code. As described in Section VI, the ability of LLMs to fix complex security vulnerabilities often requires a deep contextual understanding of the code and its dependencies, deeper than what current LLMs offer. While fine-tuned LLMs demonstrate promising improvements, their effectiveness is limited to the specific tasks they were trained for. Therefore, as it stands, LLMs need to be significantly improved for allowing them to be used as standalone tools for fixing vulnerabilities.
@@
Finally, as discussed in Subsection IX-A, the literature reveals that poisoned data can influence LLMs to generate vulnerable code. To address this challenge, future research should focus on mitigating the potential impact that malicious data could have on the security of LLM-generated code. Despite the advancements in understanding poisoning attacks on LLMs, as discussed in Subsection IX-B, there is a lack of literature work that addresses how poisoning attacks could also affect the ability of LLMs to detect and fix vulnerabilities. Given that poisoned models have been shown being capable of suggesting insecure code, it is reasonable to hypothesize that poisoning could affect LLMs ability to effectively detect and fix security vulnerabilities in code (whether computer- or human-generated).
X-C Threats to Validity
The rapid evolution of LLMs presents challenges to the validity of this research. As LLMs undergo continuous improvements and updates, their capabilities in generating code, detecting vulnerabilities, and fixing security flaws evolve.
@@ X-C Threats to Validity
Additionally, a potential threat to validity is the fact that varying prompts are used across different studies that answer RQ1. These prompts may influence the generated code, potentially resulting in code that is either more or less vulnerable.
XI Conclusion
@@
Acknowledgment
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. We would also like to express our gratitude to Markus Kuchler for his continuous support throughout the course of our research. His participation has been invaluable to the progress of this work.

### v3

@@ 9.2. Impact of Poisoned Training Data on Ability to Detect and Fix Vulnerabilities
Although research on data poisoning in LLMs exists, there are currently no studies that examine its particular effects on vulnerability detection and fixing tasks. This gap in the literature will be further discussed in the Open challenges, in Section 10.2.
10. Discussion
@@
In Section 3, as a part of our methodology, we have defined three RQs that steered the focus of this study. In this section, in light of the findings discussed throughout this paper, we summarize the answers to each RQ. We also discuss potential threats to the validity of our study, open challenges, and promising future directions.
10.1. Addressing Research Questions
@@
The literature revealed that LLM-generated code often contains different security vulnerabilities. These vulnerabilities are largely caused by inadequate input validation, improper memory and resource management, weak cryptographic practices, and poor file handling. Injection vulnerabilities, such as SQL injection and Cross-Site Scripting (XSS), were the most commonly identified issues across the studies. Although LLMs can speed up the code generation process, they can produce vulnerable code.
@@
Studies have shown that LLMs trained on poisoned datasets can generate insecure code patterns, increasing the risk that developers unknowingly introduce vulnerabilities. Additionally, studies have proposed different poisoning attack methods that can successfully bypass detection by SATs and LLMs, for example, by hiding malicious snippets in less visible places. Studies have shown that even a small fraction of poisoned training data (i.e., less than 3%) could lead a model to produce vulnerable code. Currently, no research addresses in-depth the effectiveness of poisoned LLM models, when tasked to detecting and fixing vulnerabilities.
10.2. Open Challenges
@@
Although LLMs show great potential in code generation and security-related tasks, they still face challenges that need to be addressed. As discussed in Section 4, LLM-generated code often introduces security vulnerabilities. Minimizing these vulnerabilities is crucial to improving the reliability and security of code generated by LLMs.
@@
Besides the aforementioned challenges faced in producing secure code, LLMs encounter challenges also when analyzing code for detecting vulnerabilities in it. As we have uncovered in Subsection 5.2, the first and main issue is the one of false positives, which severely limits current LLMs capabilities in vulnerability detection. Reducing false positive rates is necessary for ensuring higher reliability of LLMs and their broader adoption for undertaking these tasks. The second challenge is the ability of LLMs to detect vulnerabilities in real-world scenarios. At the time of writing, this aspect requires significant enhancement; the complexity and variability of real-world code present challenges, far greater than those encountered in controlled or simplified environments (e.g., most vulnerability datasets).
@@
Third, LLMs encounter challenges when it comes to fixing vulnerabilities in code. As described in Section 6, the ability of LLMs to fix complex security vulnerabilities often requires a deep contextual understanding of the code and its dependencies, deeper than what current LLMs offer. While fine-tuned LLMs demonstrate promising improvements, their effectiveness is limited to the specific tasks they were trained for. Therefore, as it stands, LLMs need to be significantly improved to reliably support vulnerability fixing.
@@
Finally, as discussed in Subsection 9.1, the literature reveals that poisoned data can influence LLMs to generate vulnerable code. To address this challenge, future research should focus on mitigating the potential impact that malicious data could have on the security of LLM-generated code. Despite the advancements in understanding poisoning attacks on LLMs, as discussed in Subsection 9.2, there is a lack of literature that addresses how poisoning attacks could also affect the ability of LLMs to detect and fix vulnerabilities. Given that poisoned models have been shown being capable of suggesting insecure code, it is reasonable to hypothesize that poisoning could affect LLMs ability to effectively detect and fix security vulnerabilities in code (whether computer- or human-generated).
10.3. Threats to Validity
@@
The rapid evolution of LLMs presents challenges to the validity of this research. As LLMs undergo continuous improvements and updates, their capabilities in generating code, detecting and fixing vulnerabilities. Additionally, a potential threat to validity is the fact that varying prompts are used across studies addressed RQ1. These prompts may influence the generated code, potentially resulting in code that is either more or less vulnerable.
11. Conclusion